Which function is the inverse of y=-3x+6/7?
A. y=-1/3x+2/7
B. y=-1/3x+7/6
C. y=-3x-6/7
D. y=-1/3x+18/7

Answer by rothauserc(4718)

inverse of y =-3x+6/7
solve equation for x
add -3x to both sides of the =
3x +y = 6/7
subtract y from both sides of =
3x = -y +6/7
divide both sides of = by 3
x = -y/3 +2/7
now interchange x's and y's
inverse is y = -x/3 +2/7
the answer is A.